[Brahms-l] Shift report 20040202 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 02 2004 - 08:07:55 EST
------------------------------------------------------------
Date: Mon Feb  2 08:07:54 2004, Shift: 00:00-08:00
Supervisor: rk, with: 
Log book#15 pages 48 - 48

Shift Summary:
------------------------------------------------------------
Finished 6 degrees program for FS: 
A1723 - runs 10022 - 10024 
runs 10026, 10027, 10028 were the zero-field runs. 

No problems with detectors, except of quite funny behaviour for TOFW 61 observed: fluctuations in voltages of about 15-20V. 
SuperMon has some problems just now, it prints only the first page.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 2 08:11:48 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 02 2004 - 08:12:07 EST